OpenStudy (anonymous):

Simplify sqrt 13f^3/sqrt 5f^2

OpenStudy (triciaal):



OpenStudy (anonymous):

yes, thank you!

OpenStudy (anonymous):

I started dividing out numerator and denominator by f2 leaving \[\sqrt{13f}\] /\[\sqrt{5}\]

OpenStudy (anonymous):

then rationalized the denominator by multiplying by \[\sqrt{5}\]

OpenStudy (anonymous):

leaving \[\frac{ \sqrt{13f}\times \sqrt{5} } 5\]

OpenStudy (anonymous):

Just not sure where to go from here, help anyone?

OpenStudy (anonymous):

Is \[\frac{ \sqrt{65f} }{ 5 }\] the simplest form?